OpenStudy (mayankdevnani):

A block of wood of length 40cm and area of cross section 15cm^2 floats in water with 3/8 of its length above water.What is the density of wood? (Density of water=1g/cm^3) a) 0.256g/cm^3 b)0.526g/cm^3 c)0.625g/cm^3 d)0.650g/cm^3

OpenStudy (mayankdevnani):

@UnkleRhaukus @sauravshakya @Coolsector @hartnn @.Sam. @bahrom7893

OpenStudy (mayankdevnani):

@estudier @aroub

OpenStudy (mayankdevnani):

@ganeshie8 @Yahoo!

OpenStudy (anonymous):

buoyancy force will be 15 * (5 * 40 / 8) * 1 * 9.8 upwards .. now we have gravitational force Mg we can change Mg into (densityWood) * (volumeWood) * 9.8 so Mg is (densityWood) * 40 * 15 * 9.8 now Mg and buoyancy are equal since the wood stands still : densityWood * 40 * 15 * 9.8 = 15 * 25 * 9.8 densityWood = 0.625

OpenStudy (anonymous):

sorry for all the corrections .

OpenStudy (anonymous):

you should read it all over again since i corrected stuff at the first line as well

OpenStudy (anonymous):

@Coolsector I guess buoyancy force will be 15 * (3 * 40 / 8) * 1 * 9.8 ^ Check this

OpenStudy (anonymous):

@sauravshakya with 3/8 of its length above water ABOVE WATER

OpenStudy (anonymous):

use formula v1.d1.g= v2. d2.g

OpenStudy (vincent-lyon.fr):

You do not need gravity to solve this problem. If an object sinks: volume of object = volume of (displaced) water If an object floats: mass of object = mass of (displaced) water

OpenStudy (anonymous):

obviously..but this will help dude to get insight of full fromula

OpenStudy (mayankdevnani):

@hartnn @sauravshakya @Yahoo!

OpenStudy (mayankdevnani):

i don't understand

OpenStudy (anonymous):

So, u didnt get what @coolsector did?

OpenStudy (mayankdevnani):

no

OpenStudy (mayankdevnani):

plz tell me

OpenStudy (anonymous):

Upthrust = weight of displaced liquid

OpenStudy (anonymous):

U know that?

OpenStudy (mayankdevnani):

yaa

OpenStudy (anonymous):

And when a body floats the weight of the body is equal to the upthrust

OpenStudy (anonymous):

right?

OpenStudy (mayankdevnani):

right

OpenStudy (anonymous):

Thus, weight of the object=weight of displaced liquid Mass of the object=mass of displaced liquid density*volume of object=density*volume of displaced liquid density*40*15 = 1* 5/8 *40 *15

OpenStudy (anonymous):

Now, simplify it.

OpenStudy (mayankdevnani):

375

OpenStudy (mayankdevnani):

it gets 375

OpenStudy (anonymous):

Nope it is 5/8

OpenStudy (mayankdevnani):

howz

OpenStudy (anonymous):

Are u serious??????? Just do simple algebra.

OpenStudy (mayankdevnani):

\[1\times \times \frac{5}{8}\times40\times15\]

OpenStudy (anonymous):

I know u can do it.

OpenStudy (mayankdevnani):